# Given Enough Time

### v1

@@ 4 Discussion
Our results strikingly demonstrate that the human visual system is more robust than deep networks at recognizing objects in unusual poses. Indeed, with unlimited viewing time, human performance is not affected by unusual poses, whereas deep networks are fooled by them. Only in the very limited viewing time condition (40 ms) does the performance of humans deteriorate to the level of deep networks.
@@ 4.1 The gap still exists
Regarding the deep networks, it is notable that the architecture (e.g., visual transformers vs. convolutional neural networks), loss function (e.g., self-supervised vs. supervised) and training set size (e.g., 3.6 billion images for SWAG) did not affect their performance on the task. This observation adds to a literature showing that scaling alone seems insufficient to match human internal representations and performance in visual tasks (e.g., (Shankar et al., 2020; Fel et al., 2022; Muttenthaler et al., 2022; Linsley et al., 2023a, b)). We also note that networks do not benefit from additional time, as their processing time is fixed by the nature of their feed-forward architecture.
@@ 4.2 Why do humans need the extra time?
Our results are in line with a tradition of psychology studies (Shepard & Metzler, 1971; Jolicoeur, 1985; Sofer et al., 2015; Jeurissen et al., 2016) showing that humans need more time to accomplish more challenging visual tasks. Closest to our work, Jolicoeur (1985) showed that the naming time of sketches of everyday objects was proportional to the difference in orientation with their upright pose. Below, we list some of the possible mental processes that could be responsible for this extra time:
@@
4.3 Time-limited humans are not feed-forward networks
@@
4.4 On the role of retrospective thinking in humans
In the current study, humans were presented with an image followed by a two-alternative-forced-choice question. This provides humans an opportunity to utilize retrospective thinking. For instance, a subject who initially failed to identify an image of a rotated stretcher, could reconsider their interpretation after seeing the answer alternatives ’stretcher’ and ’matchstick’. Networks, on the other hand, are not capable of such retrospective thinking; they classify the object based on the predefined 1000 classes without the possibility of retrospective thought process when given the two possible answers. Intriguingly, a recent paper of Jaini et al. (2023) proposes to use diffusion models as classifiers, by providing different possible labels to the model and using the model to retrofit the image based on these different labels. They show that such diffusion models have similar shape-biases than humans and are also capable of detecting both percepts in bistable visual illusions. However, they report that their models are still brittle when tasked to recognize rotated objects. Although they currently don’t work on this task, the idea of having models able to retrofit concepts to an image is very promising.
@@ Acknowledgements
Appendix A Deep Networks details
A.1 Dataset and Training Objectives
@@ Appendix B The 3D objects
    Figure 40: 1. Strainer, 2. Ballpoint pen, 3. Ballpoint pen
@@
  • Traffic-light: link

### v2

@@ 4 Discussion
Our results strikingly demonstrate that the human visual system is more robust than deep networks for vision and than most vision-language models at recognizing objects in unusual poses. Indeed, with unlimited viewing time, human performance is not affected by unusual poses, whereas deep networks are fooled by them. Only in the very limited viewing time condition (40 ms) does the performance of humans deteriorate to the level of deep networks.
@@ 4.1 The gap still exists
Regarding the deep networks, it is notable that the architecture (e.g., visual transformers vs. convolutional neural networks), loss function (e.g., self-supervised vs. supervised), training set size (e.g., 3.6 billion images for SWAG) and modality (pure images, or images and text) did not affect their performance on the task (with the exception of Gemini, see discussion below). This observation adds to a literature showing that scaling alone seems insufficient to match human internal representations and performance in visual tasks (e.g., (Shankar et al., 2020; Fel et al., 2022; Muttenthaler et al., 2022; Linsley et al., 2023a, b)). We also note that networks do not benefit from additional time, as their processing time is fixed by the nature of their feed-forward architecture.
@@ 4.2 Why do humans need the extra time?
Our results are in line with a tradition of psychology studies (Shepard & Metzler, 1971; Jolicoeur, 1985; Edelman & Bülthoff, 1992; Kosslyn et al., 1994; Sofer et al., 2015; Jeurissen et al., 2016; Kallmayer et al., 2023; Mayo et al., 2024) showing that humans need more time to accomplish more challenging visual tasks. Closest to our work, Jolicoeur (1985) showed that the naming time of sketches of everyday objects was proportional to the difference in orientation with their upright pose. Below, we list some of the possible mental processes that could be responsible for this extra time:
@@
4.3 Time-limited humans are not feed-forward deep networks
@@
4.4 On the role of retrospective thinking in humans and vision-language models
In the current study, humans were presented with an image followed by a two-alternative-forced-choice question. This provides humans with an opportunity to utilize retrospective thinking. For instance, a subject who initially failed to identify an image of a rotated stretcher, could reconsider their interpretation after seeing the answer alternatives ’stretcher’ and ’matchstick’. Deep networks for pure vision, on the other hand, are not capable of such retrospective thinking; they classify the object based on the predefined 1000 classes without the possibility of a retrospective thought process when given the two possible answers. Conversely, large vision-language models are given the opportunity to ponder the content of the image after being fed the two labels, like humans. Their performance is notably better than deep networks for pure vision (except Claude), and even matching human performance in the case of Gemini. The impressive performance of vision-language models suggests that they might be able to exploit similar mechanisms for image recognition as humans, which could for instance involve retro-fitting the labels to the image. However, an alternative explanation is that they could have been fed a different data diet involving more rotated objects than their pure vision counterparts. Because of the limited documentation available for these models, it is difficult to come to a firm conclusion regarding the mechanisms underlying their abilities.
@@ Acknowledgements
Appendix A Testing Large Vision-Language Models
We tested 6 large vision-language models, Gemini 1.5 Flash, Gemini 1.5 Pro, GPT-4o, GPT-4-vision-preview, Claude 3.5 Sonnet and Claude 3 Opus, via their API. Figure 6 shows an example of the used prompt and models’ answers.
Figure 6: Examples of a prompt and models’ answers in API mode. a) GPT-4o, b) GPT-4-vision-preview, c) Claude 3.5 Sonnet, d) Claude 3 Opus, e) Gemini 1.5 Flash and f) Gemini 1.5 Pro (example of safety-blocked instance, which was disregarded in our analysis).
Appendix B Deep Networks (pure vision) details
B.1 Dataset and Training Objectives
@@ Appendix C The 3D objects
    Figure 41: 1. Strainer, 2. Ballpoint pen, 3. Mop
@@
  • Traffic light: link